Presented below is information related to Stellar Corporation for the current year.

Beginning inventory $ 596,100
Purchases 1,522,700
Total goods available for sale $2,118,800
Sales revenue 2,590,000


Compute the ending inventory, assuming that (a) gross profit is 45% of sales, (b) gross profit is 60% of cost, (c) gross profit is 37% of sales, and (d) gross profit is 25% of cost. (Round ratios for computational purposes to 1 decimal place, e.g. 78.7% and final answers to 0 decimal places, e.g. 28,987.)

Ending Inventory

(a) Gross profit is 45% of sales

$

(b) Gross profit is 60% of cost

$

(c) Gross profit is 37% of sales

$

(d) Gross profit is 25% of cost

$
Answer #1
Ending Inventory
a) Gross profit is 45% of sale $694,300
b) Gross profit is 60% of cost $500,500
c) Gross profit is 37% of sales $487,100
d) Gross profit is 25% of cost $46,800
Working Note:-
a) Gross profit is 45% of sale
Cost of goods sold = Sale - Gross Profit
Cost of goods sold =2,590,000 - 45%*2,590,000
Cost of goods sold = 1,424,500
Ending Inventory = Total goods available for sale -cost of goods sold
Ending Inventory = 2,118,800 - 1,424,500
Ending Inventory = 694,300
b) Gross profit is 60% of cost
Cost of goods sold = Sale - Gross Profit
Cost of goods sold =2,590,000 - (60/160)*2,590,000
Cost of goods sold = 1,618,750
Ending Inventory = Total goods available for sale -cost of goods sold
Ending Inventory = 2,118,800 - 1,618,750
Ending Inventory = 500,500
c) Gross profit is 37% of sale
Cost of goods sold = Sale - Gross Profit
Cost of goods sold =2,590,000 - 37%*2,590,000
Cost of goods sold = 1,631,700
Ending Inventory = Total goods available for sale -cost of goods sold
Ending Inventory = 2,118,800 - 1,631,700
Ending Inventory = 487,100
d) Gross profit is 25% of cost
Cost of goods sold = Sale - Gross Profit
Cost of goods sold =2,590,000 - (25/125)*2,590,000
Cost of goods sold = 2,072,000
Ending Inventory = Total goods available for sale -cost of goods sold
Ending Inventory = 2,118,800 - 2,072,000
Ending Inventory = 46,800